# From First Patch to Long-Term Contributor: Evaluating Onboarding Recommendations for OSS Newcomers

### v1

@@ 3.2.1 Recommendation to quantitative metrics mapping
We select a quantitative metric for each recommendation based on prior works on code reviews [59, 61, 62, 63, 64], and the 14 newcomers’ onboarding-related studies found in our SLR. Identifying metrics for R2, R4, R5, and R8-R11 was straightforward, as those directly allude to measures. While task size (R1) can be measured using various metrics, we use ‘total churn’, the total number of lines added, modified, or deleted in a change, since this metric has been used by prior code review studies to estimate change size [59, 62]. In addition to new feature-related labels, we map adding new functionality (R3) to whether a new file was added since a significant share of pull requests do not have any label. We compute the ‘Change entropy’ metrics introduced by Hassan [58] to measure task complexity (R6), as this metric has also been used in several code review-related studies to compute change complexity [59, 62]. To map R7: working on documentation, we compute the ratio of files in the change with documentation-related extensions. To map R12: readable titles and R13: descriptions, we select Flesch Reading ease [60], since it is one of the most widely used readability scores and prior research has found this metric suitable to assess readabilities of technical manuals [65]. We map R14: authors’ proactivity with author promptness (i.e., how quickly the author resubmitted after a review) and R15: being patient with the number of iterations (i.e., how many times the author resubmitted the change to address suggested changes). Table III shows the mapping among the 15 recommendations and their attribute mappings.
@@ 3.2.2 Sample Selection
For this study, we selected projects from two sources. First, we selected five projects using Gerrit-based code reviews to accept newcomers’ code changes. Gerrit is among the most popular code review tools many non-GitHub-based OSS projects use. Our selection criteria for Gerrit projects include i) Has been using Gerrit for at least five years, ii) Has at least 50,000 CRs, and iii) Has at least 500 newcomers. Based on these criteria, we identified the following five from 21 Gerrit-based OSS projects [66]: i) Android: a mobile operating system sponsored by Google; ii) LibreOffice: an office development platform sponsored by a nonprofit foundation; iii) Typo3: a web content management system run by a nonprofit community; iv) oVirt: a platform for managing virtual machines, managed by a for-profit company; and v) Wikimedia: wiki management software, community-based. These five projects represent diverse domains, sponsorship, and programming languages. These projects have also been subject to prior code review studies [67, 63]. Next, we selected OSS projects using the PR-based development model [68] to accept newcomers’ code changes. We use a GitHub project search tool developed by Dabic et al. [69] for this goal. Following the recommendations of Kalliamvakou et al. [70], we excluded projects that do not meet the following four criteria.
@@ 3.2.3 Dataset Preparation
Table IV shows an overview of our dataset. Gerrit data: We have used a Gerrit Miner tool [72] to collect all the submitted review requests till 30th April, 2023. Our Gerrit dataset includes 1,438,042 code review (CR) requests from the five selected projects. We adopted a keyword-based filtering approach proposed by Bosu et al. [67] to identify and exclude bot-generated interactions. With this approach, accounts with the following set of keywords were marked: ‘bot’, ‘CI’, ‘Jenkins’, ‘build’, ‘auto’, ‘devop’, ‘hook’, ‘workflow’, and ‘travis’. Next, we manually inspected each marked account to make a final determination. We leverage the Active Learning Fingerprint Based Anti-Aliasing (ALFAA) approach proposed by Amreen et al. [73] to identify multiple Gerrit accounts belonging to the same person into one account.
@@
Table IV: Overview of the selected dataset
@@ 3.2.4 Attribute Calculation
We wrote Python scripts and MySQL queries to compute the metrics listed in Table III. However, we had to define additional heuristics for the following set of metrics.
  • Is newcomer: Following the recommendation of Steinmacher et al., we consider a contributor as a newcomer to a project till he/ she/ they have got their first CR /PR accepted [5, 13].
@@ 3.2.5 Model Training and Evaluation
We train Multivariate Logistic Regression (MLR) models to identify associations between attributes listed in Table III and CR/PR acceptance. Although MLR has been recently used in Machine learning, it has been used for decades by statisticians since its invention by David Cox in 1958 [76]. MLR effectively investigates the relationships between a binary dependent variable and one or multiple independent variables [77]. There are two categories of MLR models: i) predictive – where the goal is to predict future outcomes based on past data, and ii) inferential – where the goal is to identify associations between the dependent and multiple independents. Predictive models are evaluated using precision, recall, F-score, and accuracy. On the other hand, inferential models are evaluated using their ability to explain observed variances, usually measured using Pseudo-$R^{2}$  [18]. Based on our study’s goal, we train inferential MLR models.
@@
Multicollinearity arises when two or more independent variables are highly correlated and is a threat to inferential MLR models [78]. We followed the approach of Harell Jr. [18], which has been used in many prior SE studies [59, 61, 64]. Our models use the 15 factors listed in Table III as the independents and acceptance of PR/CR as the dependent variable. We used 14 independent variables for Gerrit projects, as GFI tags are not used among Gerrit-based projects. We named our dependent variable as is_accepted, set to ‘1’ if the patch is accepted and ‘0’ otherwise. While computing is_accepted was straightforward for Gerrit projects, this variable may be incorrect on Github [68, 79]. We followed the heuristics suggested by Gousios et al. [68] to resolve potentially incorrect acceptance status on GitHub. For each of the eight datasets (i.e., 5 Gerrit and 3 GitHub), we trained two MLR models. The first model is trained using the CRs / PRs submitted by newcomers to assess RQ2. The second model is trained using non-newcomers’ PRs /CRs to answer RQ3. We log-transformed four variables (i.e., insert, delete, delay, and churn) as those had highly skewed distributions. Our model training and evaluation steps are as follows.
@@
Step 1: Correlation and Redundancy Analysis: To identify highly correlated explanatory variables, we applied Sarle’s Variable Clustering (VURCLUS) approach [80] where we used Spearman’s rank-order correlation test [81] to build a hierarchical cluster representation of explanatory variables. Following the suggestion of Hinkle et al. [82], we used the correlation coefficient $|\rho|=0.7$ as the cutoff threshold. If multiple explanatory variables belong to the same cluster with $|\rho|\geq 0.7$, we picked only one from that cluster that has the highest correlation with the response variable. Even if not highly correlated, explanatory variables can still be redundant, eventually affecting the model’s performance. We used the Autospearman R library created by Jiarpakdee et al. [83] to select the best set of factors for each model after assessing correlation and redundancy.
@@
Power of each variable: We conducted an ablation study to understand how the inclusion/exclusion of each independent influences a model’s association with the dependent. In this design, we start with a null model and compute its unexplained variance (i.e., Null deviance ($DIV_{null}$)) [^1] Deviance is a measure to assess goodness of fit for regression models. It is computed based on the difference between trained and perfect models. The lower the deviance, the better the model.. Then, we measure the change in residual deviance ($\Delta DIV_{residual}$) with the addition of each independent. Therefore, the explanatory power of an independent is computed according to the following equation: $EP$ = $\frac{\Delta DIV_{residual}}{DIV_{null}}*100$. The higher the value of EP(%), the better that attribute explains the dependent variable.
4 Results
The following subsections detail the results of our three research questions introduced in Section 1.
Table V: Research methodology for studies on newcomers’ recommendations
Research method # Number of studies
Interview / survey 9 SLR1, SLR2, SLR4, SLR6, SLR7, SLR8, SLR9, SLR10, SLR11
Mining software repositories (MSR) 4 SLR3, SLR12, SLR13, SLR14
Mixed: Survey and MSR 1 SLR5
Table VI: Task-related recommendations for prospective newcomers identified based on our Systematic Literature Review
#ID Recommendation Rationale Studies
Recommendation group: Task selection
R1 Work on smaller tasks first and later choose larger tasks Working on smaller tasks might increase the likelihood of newcomers’ first patch getting accepted SLR1[27], SLR2[14], SLR3[39], SLR5[15], SLR6[26], SLR8[33], SLR11[12], SLR12[13]
R4 Start your contribution by working on minor bug fixes Minor bug fixes have lower knowledge requirements and also help improve the quality of a project SLR3[39], SLR5[15], SLR12[13], SLR14[32]
R6 Start your contribution by performing less complex tasks Complex tasks are difficult for newcomers to solve and can possibly demotivate newcomers due to rejections SLR1[27], SLR3[39], SLR4[20], SLR7[30], SLR8[33], SLR9[86], SLR10[5], SLR13[10], SLR14[32]
R7 Start your contribution by working on documentation changes Documentation changes are less complex, have fewer contribution barriers such as setting up environment or testing, and therefore have a higher odds of getting accepted SLR3[39], SLR5[15], SLR12[13], SLR14[32]
Recommendation group: Task preparation
Recommendation group: Task completion
Table V shows an overview of the research method from the 14 selected primary studies. Nine out of the 14 studies are qualitative, and most of those recommendations did not include any quantitative evaluation of those recommendations. The four quantitative studies are recent works on characterizing GFIs on GitHub but lack analysis of non-GFIs. This result further shows the need for a large-scale quantitative analysis. Table VI shows our list of 15 recommendations identified based on our SLR. We divided the recommendations into three groups: Task selection, Task preparation, and Task completion. Under the task selection group, we got nine recommendations. The two most popular recommendations within this group for newcomers are to work on less complex tasks (R6), found in 9 primary studies, and to work on smaller tasks (R1), found in 8 studies. Within the Task preparation group, we obtained four recommendations, where making the title and description comprehensible ranks first with three studies. We found two Task completion recommendations (i.e., what to do after patch submission). While most recommendations are congruent, we found a conflict between R3 and R4, i.e., starting with new features vs. bug fixes.
@@ 4.1 RQ1: What are the task-related recommendations for prospective newcomers of OSS projects?
For a project, the $OR^{p}$ columns in Tables VII and VIII list effect sizes for the attributes, estimated using Odds ratio. For example, in the LibreOffice project, the $OR^{p}$ value for the newcomer group for the doc file factor is 4.56 (Table VII). This result indicates that if a newcomer’s change comprises 100% doc files in LibreOffice, the odds of acceptance are 4.56 times higher than that of newcomers’ changes, including 0% doc files. The EP(%) columns in Tables VII and VIII indicate the explanatory power of each attribute for the model, with higher EP(%)s indicating better predictors. For example, in Table VII, the value of EP(%) for bugfix is 4.47% in Android indicating percentage contribution of that variable for Android’s model. We consider a recommendation supported only if the mapped factor has significant positive associations for a project. In case of negative associations, the recommendation is contradicted. The recommendation is insignificant if no significant association ($p<0.05$) is seen.
@@
R2: GFI, which is only applicable for GitHub datasets, has significant positive associations for GitHub (M) and GitHub (H) but is insignificant for GitHub (L). R6: Starting with low complexity changes, it is supported across four datasets and insignificant for four Gerrit projects except for oVirt. R7: Working on documentation is supported across five of the eight datasets and is one of the predictors with higher explanatory powers in GitHub. We noticed context-dependencies for R3, R9, R14, and R15. While R3: adding new features significantly lowers acceptance odds on GitHub (M) and GitHub (H), an opposite association is seen in Android. Similarly, contradicting R9: fewer lines deletions222interpretations for R6, R8, and R9 are opposite of the associations with mapped variables since those suggest changes with lower complexities or fewer line insertions /deletions, but the mapped variables measure complexity and the number of lines inserted/deleted, the odds of acceptance increase with the number of lines deleted in Android and oVirt. However, we notice significant support for R9 across all three GitHub datasets. R14: being practive is supported for four datasets (i.e., LibreOffice, Typo3, Wikimedia, and Github (M)), contradicts for two (Android and Github (H)), and insignificant for the remaining two. R15: being patient is supported across all Gerrit projects and GitHub (M) but contradicted in GitHub (H). Task completion-related recommendations (i.e., R14 and R15) have high explanatory powers across all datasets and are important predictors.
@@ Acknowledgment
Work conducted for this research is partially supported by the US National Science Foundation under Grant No. 1850475. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.

### v2

@@ 4.1.1 Recommendation to quantitative metrics mapping
We select a quantitative metric for each recommendation based on prior works on code reviews [67, 69, 70, 71, 72], and the 16 newcomers’ onboarding-related studies found in our SLR. Identifying metrics for R2, R4, R5, and R8-R11 was straightforward, as those directly allude to measures. While task size (R1) can be measured using various metrics, we use ‘total churn’, the total number of lines added, modified, or deleted in a change, since this metric has been used by prior code review studies to estimate change size [67, 70]. In addition to new feature-related labels, we map adding new functionality (R3) to whether a new file was added since a significant share of pull requests do not have any label. We compute the ‘Change entropy’ metrics introduced by Hassan [66] to measure task complexity (R6), as this metric has also been used in several code review-related studies to compute change complexity [67, 70]. To map R7: working on documentation, we compute the ratio of files in the change with documentation-related extensions. To map R12: readable titles and R13: descriptions, we select Flesch Reading ease [68], since it is one of the most widely used readability scores and prior research has found this metric suitable to assess readabilities of technical manuals [73]. We map R14: authors’ proactivity with author promptness (i.e., how quickly the author resubmitted after a review) and R15: being patient with the number of iterations (i.e., how many times the author resubmitted the change to address suggested changes). Table V shows the mapping among the 15 recommendations and the selected metrics.
@@ 4.1.2 Sample Selection
For this study, we selected projects from two sources. First, we selected five projects using Gerrit-based code reviews to accept newcomers’ code changes. Gerrit is among the most popular code review tools many non-GitHub-based OSS projects use. Our selection criteria for Gerrit projects include i) Has been using Gerrit for at least five years, ii) Has at least 50,000 CRs, and iii) Has at least 500 newcomers. We selected these five criteria to ensure adequate newcomers’ CRs, which is a requisite to train reliable regression models. We identified a total of 21 publicly available Gerrit-managed CR repositories [74], where six satisfied our selection criteria. Between the two Google-managed projects, we selected Android instead of Chromium OS since it had a higher number of newcomers during our study period. We selected all four non-Google projects satisfying our criteria. Therefore, our selection of Gerrit-based OSS projects are i) Android: a mobile operating system sponsored by Google; ii) LibreOffice: an office development platform sponsored by a nonprofit foundation; iii) Typo3: a web content management system run by a nonprofit community; iv) oVirt: a platform for managing virtual machines, managed by a for-profit company; and v) Wikimedia: wiki management software, community-based. These five projects represent diverse domains, sponsorship, and programming languages. These projects have also been subject to prior code review studies [75, 71]. Next, we selected OSS projects using the PR-based development model [76] to accept newcomers’ code changes. We use a GitHub project search tool developed by Dabic et al. [77] for this goal. Following the recommendations of Kalliamvakou et al. [78], we excluded projects that do not meet the following four criteria.
@@ 4.1.3 Dataset Preparation
Table VI shows an overview of our dataset. Gerrit data: We have used a Gerrit Miner tool [80] to collect all the submitted review requests till 30th April, 2023. Our Gerrit dataset includes 1,438,042 code review (CR) requests from the five selected projects. We adopted a keyword-based filtering approach proposed by Bosu et al. [75] to identify and exclude bot-generated interactions. With this approach, accounts with the following set of keywords were marked: ‘bot’, ‘CI’, ‘Jenkins’, ‘build’, ‘auto’, ‘devop’, ‘hook’, ‘workflow’, and ‘travis’. Next, we manually inspected each marked account to make a final determination. We leverage the Active Learning Fingerprint Based Anti-Aliasing (ALFAA) approach proposed by Amreen et al. [81] to identify multiple Gerrit accounts belonging to the same person into one account.
@@
Table VI: Overview of the selected dataset
@@ 4.1.4 Attribute Calculation
We wrote Python scripts and MySQL queries to compute the metrics listed in Table V. However, we had to define additional heuristics for the following set of metrics.
  • Is newcomer: Following the recommendation of Steinmacher et al., we consider a contributor as a newcomer to a project till he/ she/ they have got their first CR /PR accepted [5, 13]. We consider all other contributors as non-newcomers and use them as the ‘control group’. However, we acknowledge that non-newcomers are diverse and can be further divided into several sub-categories [25, 21, 28, 29].
@@ 4.1.5 Model Training and Evaluation
We train Multivariate Logistic Regression (MLR) models to identify associations between attributes listed in Table V and CR/PR acceptance. Although MLR has been recently used in Machine learning, it has been used for decades by statisticians since its invention by David Cox in 1958 [84]. MLR effectively investigates the relationships between a binary dependent variable and one or multiple independents [85]. There are two categories of MLR models: i) predictive – where the goal is to predict future outcomes based on past data, and ii) inferential – where the goal is to identify associations between the dependent and multiple independents. Predictive models are evaluated using precision, recall, F-score, and accuracy. On the other hand, inferential models are evaluated using their ability to explain observed variances, usually measured using Pseudo-$R^{2}$  [18]. Based on our study’s goal, we train inferential MLR models.
@@
Multicollinearity arises when two or more independent variables are highly correlated and is a threat to inferential MLR models [86]. We followed the approach of Harell Jr. [18], which has been used in many prior SE studies [67, 69, 72]. Our models use the 15 factors listed in Table V as the independent variables and acceptance of PR/CR as the dependent variable. We used 14 independents for Gerrit projects, as GFI tags are not used among Gerrit-based projects. We named our dependent variable as is_accepted, set to ‘1’ if the patch is accepted and ‘0’ otherwise. While computing is_accepted was straightforward for Gerrit projects, this variable may be incorrect on Github [76, 87]. We followed the heuristics suggested by Gousios et al. [76] to resolve potentially incorrect acceptance status on GitHub. For each of the eight datasets (i.e., 5 Gerrit and 3 GitHub), we trained two MLR models. The first model is trained using the CRs / PRs submitted by newcomers to assess RQ2. The second model is trained using non-newcomers’ PRs /CRs to answer RQ3. We log-transformed four variables (i.e., insert, delete, delay, and churn) as those had highly skewed distributions. Our model training and evaluation steps are as follows.
@@
Step 1: Correlation and Redundancy Analysis: To identify highly correlated explanatory variables (i.e., Multicollinearity), we applied Sarle’s Variable Clustering (VURCLUS) approach [88]. We used Spearman’s rank-order correlation test [89] to build a hierarchical cluster representation of explanatory variables. Following the suggestion of Hinkle et al. [90], we used the correlation coefficient $|\rho|=0.7$ as the cutoff threshold. If multiple explanatory variables belong to the same cluster, we picked only the one with the highest correlation with the dependent. For example, in most models, we found that log(code churn) and log (insert) are highly correlated (i.e., $|\rho|\geq 0.7$). Since log(insert) has a higher correlation with the dependent, we dropped log(code churn). Even if not highly correlated, explanatory variables can still be redundant, eventually affecting the model’s performance. We use the redun function of the rms R package with the threshold $R^{2}\geq$ 0.9 to identify potential redundant factors among the ones surviving the multicollinearity threshold.
@@
Power of each variable: We conducted an ablation study to understand how the inclusion/exclusion of each independent variable influences a model’s association with the dependent. In this design, we start with a null model and compute its unexplained variance (i.e., Null deviance ($DIV_{null}$)) [^2] Deviance is a measure to assess goodness of fit for regression models. It is computed based on the difference between trained and perfect models. The lower the deviance, the better the model.. Then, we measure the change in residual deviance ($\Delta DIV_{residual}$) with the addition of each independent. Therefore, the explanatory power of an independent is computed according to the following equation: $EP$ = $\frac{\Delta DIV_{residual}}{DIV_{null}}*100$. The higher the value of EP(%), the better that attribute explains the dependent variable.
@@
For a project, the $OR^{p}$ columns in Tables VII and VIII list effect sizes for the attributes, estimated using Odds ratio. For example, in the LibreOffice project, the $OR^{p}$ value for the newcomer group for the doc file factor is 4.56 (Table VII). This result indicates that if a newcomer’s change comprises 100% doc files in LibreOffice, the odds of acceptance are 4.56 times higher than that of newcomers’ changes including 0% doc files, while holding the other variables constant. The EP(%) columns in Tables VII and VIII indicate the explanatory power of each attribute for the model, with higher EP(%)s indicating better predictors. For example, in Table VII, the value of EP(%) for bugfix is 4.47% in Android indicating percentage contribution of that variable for Android’s model. We consider a recommendation supported only if the mapped factor has significant positive associations for a project. In case of negative associations, the recommendation is contradicted. The recommendation is insignificant if no significant association ($p<0.05$) is seen.
@@
R2: GFI, which is only applicable for GitHub datasets, has significant positive associations for GitHub (M) and GitHub (H) but is insignificant for GitHub (L). R6: Starting with low complexity changes, it is supported across four datasets and insignificant for four Gerrit projects except for oVirt. R7: Working on documentation is supported across five of the eight datasets and is one of the predictors with higher explanatory powers in GitHub. We noticed context-dependencies for R3, R9, R14, and R15. While R3: adding new features significantly lowers acceptance odds on GitHub (M) and GitHub (H), an opposite association is seen in Android. Similarly, contradicting R9: fewer lines deletions [^3] Interpretations for R6, R8, and R9 are opposite of the associations with mapped variables since those suggest changes with lower complexities or fewer line insertions /deletions, but the mapped variables measure complexity and the number of lines inserted/deleted., the odds of acceptance increase with the number of lines deleted in Android and oVirt. However, we notice significant support for R9 across all three GitHub datasets. R14: being practive is supported for four datasets (i.e., LibreOffice, Typo3, Wikimedia, and Github (M)), contradicts for two (Android and Github (H)), and insignificant for the remaining two. R15: being patient is supported across all Gerrit projects and GitHub (M) but contradicted in GitHub (H). Task completion-related recommendations (i.e., R14 and R15) have high explanatory powers across all datasets and are important predictors.
@@ 5.3 Implications
Future directions: The results of our study found three newcomer-specific recommendations related to task selection, suggesting that an OSS contributor should change his/her task selection strategies after his/her first patches are accepted in an OSS project if he/she intends to become a long-term contributor to that project. However, non-newcomers groups are diverse, as prior OSS literature has further categorized non-newcomers into various groups such as quasi-contributors [27], casual contributors [21], episodic volunteers [29], core and peripheral developers. Based on our results, it is plausible that task-related strategies may further differ among various non-newcomer groups. Such investigations will allow customizing recommendations for a non-newcomer based on his/her individual goals. We consider those analyses as one of the future directions for this research.
@@ 6 Threats to Validity
Combining various categories of non-newcomers into one group to train one regression model poses a threat to conclusion validity, since non-newcomers are diverse [93, 29, 1, 21]. Hence, some recommendations may apply differently to a non-newcomer subgroup than the whole. A fine-grained investigation is essential to identify those nuanced differences.
@@ Acknowledgment
Work conducted for this research is partially supported by the US National Science Foundation under Grants 1850475 and 2340389. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.